# Efficiently serving large multimedia models using EPD Disaggregation

### v1

@@ 5.3 Memory Savings through Stage Disaggregation
Table 3: Comparison of maximum supported E and P batch size. As seen, E, P batch sizes is much higher in EPD system in comparison to DistServe. Number of images/ request is fixed to 10. KV cache is allocated as 80% of the available free memory. Experiment done on one A800 GPU with 82GB memory. Higher (italicized) is better.
EPD Supports Larger KV Cache Sizes: Tab. 4 compares the maximum KV cache size that can be allocated in the disaggregated (EPD) system with that of the aggregated (DistServe or vLLM) system. In this experiment, the batch size is fixed at 1, and the image resolution is set to 4032×3024. We evaluate two LMM models, InternVL and MiniCPMv, with varying numbers of images per request.
The results clearly demonstrate that EPD supports significantly larger KV cache sizes than DistServe. For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for Distserve, representing a 2.2x improvement.
Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an Out of Memory (OOM) error, indicating that it cannot process 20 images even with a KV cache size of 0. Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.
@@ 40 OOM OOCL
Table 4: Comparison of maximum supported KV cache size (in terms of percentage of free memory) on prefill node for various #images/ request. Image resolution fixed to 4K. OOM refers to Out of Memory and OOCL refers to Out of Context Limit. Experiments are done on A800 GPU with 82GB memory. Higher (italicized) is better.
Figure 6: Left: The impact of tuning the number of encoding workers in the EPD method. The annotation ‘xEyP‘ indicates that x encoders and y prefill workers are allocated for EPD, and ‘7P‘ means that the DistServe always allocates 7 P workers for the encoding and prefill stage requests. Middle: The impact of the number of images per request in the workload. Right: The impact of encoding and prefill batch sizes in the EPD method (encoding and prefill batch sizes are set to be equal).
5.4 Throughput in Offline Settings
In this section, we demonstrate the benefits of the EPD method in terms of throughput, specifically focusing on its memory efficiency and the ability to support higher batch sizes for both the encoding (E) and prefill (P) stages.
Consider an offline scenario where a batch of requests is submitted, allowing the system to process them overnight with a focus on maximizing end-to-end (E2E) throughput. In the traditional DistServe method, memory is allocated for both the encoding and prefill stages on the same worker, which is constrained by the total memory capacity of the GPU.
This memory limitation becomes especially significant in heterogeneous environments with low-end GPUs that have limited memory. In these cases, the DistServe method struggles to fully utilize low-end GPUs, as the combined memory demands of the encoding and prefill stages exceed the available memory. In contrast, our disaggregated EPD method is more cost-effective, as it assigns the low-end GPUs exclusively to the encoding stage, allowing better utilization of available memory.
In this experiment, we consider a less extreme scenario where the memory demands of the prefill workers in the DistServe method do not exceed the memory capacity of the low-end GPUs, allowing for a batch size of 1 for both the encoding and prefill stages.
We use 8 A800 GPUs for this experiment. For EPD, we allocate 5 GPUs for encoding (E), 2 GPUs for prefill (P), and 1 GPU for decoding (D), with the maximum batch sizes set to 8, 8, and 128 for E, P, and D, respectively. In the DistServe method, we allocate 7 GPUs for prefill (P) and 1 GPU for decoding (D), with the maximum batch sizes of 1 and 128, respectively. The workload consists of 1,000 requests, each with a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens.
We tune the system’s hyperparameters as outlined in Sec. 3.3.3 and evaluate their impact on throughput. The results are presented in Fig. 6.
The left plot illustrates the importance of selecting an appropriate GPU configuration. As described in Sec. 3.3.3, our algorithm automatically identifies that a 5E2P configuration (5 encoder and 2 prefill workers) maximizes E2E throughput.
The middle plot shows that EPD achieves superior throughput when the number of images per request is low, demonstrating that EPD’s architecture prevents the encoder from becoming a compute-bound bottleneck.
The right plot highlights that EPD’s performance is relatively insensitive to encoding and decoding batch sizes, consistently outperforming the DistServe method. This flexibility allows us to either choose a relatively large batch size or rely on the algorithm in Sec. 3.3.3 to automatically select an optimal batch size.
6 Conclusion
Our results demonstrate significant improvements in both throughput and memory efficiency, particularly in heterogeneous environments where resources are limited. By efficiently parallelizing the encoding and prefill stages, we are able to mitigate common bottlenecks and improve the scalability of multimodal inference tasks. Furthermore, the optimized memory management strategies ensure that the system can scale to handle larger models and more complex multimodal inputs, which are increasingly common in real-world applications.
The approach presented in this work represents a significant step forward in the development of scalable and efficient systems for multimodal AI, offering a foundation for future research and practical applications in areas such as natural language processing, computer vision, and cross-modal understanding. In summary, our system provides a robust solution for improving the performance of LMMs, making them more efficient and adaptable for deployment in real-world, resource-constrained environments.

### v2

@@ 4.3 Memory Savings through Stage Disaggregation
Table 2: Comparison of maximum supported E and P batch size. #images/ request is fixed to 10. KV cache is allocated as 80% of the available free memory. Higher (italicized) is better.
4.4 Ablation Study
In this section, we analyze the impact of various components of the proposed system.
#I/R=2 #I/R=4 #I/R=6 #I/R=8
EPD 0.92 1.02 1.14 1.74
w/o IRP 1.46 (1.6x) 2.47 (2.4x) 3.37 (2.9x) 4.27 (2.5x)
Table 3: Effect of ablating IRP feature from the proposed system on TTFT (s). Disabling IRP negatively affects the TTFT (up to 2.9x worse) for various multiple images/ request (#I/R). Results are averaged over 100 requests.
Effect of IRP: We analyze the effect of ablating IRP, presented in Section 3.2.2. The experimental settings are same as TTFT experiment in Sec. 4.2. As shown in Tab. 3, removing the IRP feature negatively affects TTFT across various images/ request. Moreover, the degradation exacerbates as the number of images/ request increases. This is because the IRP feature allows parallelization of encoding load within the same request across multiple GPUs.
Goodput (r/s) $\uparrow$ TTFT (s) $\downarrow$ TPOT (s) $\downarrow$
EDP 1.25 2.12 0.031
w/o Opt. 0.56 (2.2x) 4.48 (2.1x) 0.025 (0.8x)
Table 4: Effect of Offline Optimizer. When the configuration is set randomly, the goodput is 2.2× worse on average. $\downarrow$ indicates that lower is better and $\uparrow$ means that larger is better.
Effect of Offline Optimizer: By default, the EPD system collects workload samples and finds the optimal config offline using the optimizer. To demonstrate its effect, we conduct an experiment without using the optimizer and set configs randomly. Specifically, we uniformly sample 10 configs at random and report the estimated expectation of the performance metric in the second row of Tab. 4. For a fair comparison, when evaluating the TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which is the goodput of the EPD system. (See Appendix B.4 for more details) As shown in Tab. 4, disabling the optimizer causes significant performance degradation, in terms of both goodput and TTFT. This demonstrates the necessity of the optimizer, especially when the EPD system has many tunable configs.
Effect of Dynamic Role Switching: To analyze the impact of dynamic role switching, we conduct a controlled experiment simulating a shift in workload characteristics. Specifically, we generate 100 requests following the same static workload configuration as described in Section 3.2.4. However, to introduce an artificial workload change, we set the first 10 requests to generate 50 output tokens, while the remaining 90 requests generate 500 output tokens. The request arrival rate is fixed at 3 requests per second. The results are shown in Tab. 5.
Latency (s) $\downarrow$ TTFT (s) $\downarrow$ TPOT (s) $\downarrow$
EPD 28.01 1.42 0.05
w/o Switch 61.10 (2.2x) 1.33 (0.9x) 0.12 (2.4x)
Table 5: Effect of ablating dynamic role switching feature from EPD. Disabling the worker migration feature negatively affects the TPOT (2.4x worse), which results in significantly poor end-to-end latency (2.2x worse). Results are averaged over 100 requests with one 4K image per request. $\downarrow$ indicates that lower is better.
Without dynamic worker migration, the system performs poorly because it remains locked in the initial configuration (5E1P2D, optimized offline for 50 tokens) and is unable to adapt to the increased decoding demand. In contrast, the EPD system with migration dynamically reconfigures itself (2E1P5D) to handle the new workload (500 output tokens) by shifting 3 E instances to D, resulting in approximately 2× better performance.
4.5 Adaptation to Neural Processing Units (NPUs)
This section presents the results of adapting the proposed EPD framework to Huawei Ascend NPUs. For more details on the implementation and experiments, see Sec. B.5.
First, we compare the EPD-NPU against the vLLM and DistServe baselines in terms of SLO attainment using the InternVL2-8B LMM. We used the same settings as mentioned in Sec. 4.1, except that we used a heavy encoding workload of 8 4032x3024 images per request. The optimal configuration for this workload was found to be 5E2P1D and the corresponding results are presented in Figure 8. As seen, EPD is the only setting that can achieve the SLO requirements, while the other baselines are completely decimated even for low request rates.
Figure 8: SLO curve for InternVL2-8B with the following requirements (TTFT=8.5 and TPOT=0.12) on the controlled experimental setup with 8 images per request.
Next, we compare the TTFT of EPD-NPU against vLLM. We found that EPD-NPU demonstrated a 35.2% improvement over vLLM, which is significantly more than the 24.4% improvement for the GPU scenario. We attribute this additional $\sim$10to the proportionally heavier encode of LLMs on NPU compared to GPU. Please refer to Sec. D.3 for more details. This results in greater benefits of disaggregating the encode from prefill on NPU in comparison to GPU.
5 Conclusion
Appendix A Optimizer Details
We recapitulate the configuration optimization problem:
$\max_{(\mathbf{p},\mathbf{b},\mathbf{s})\in\mathcal{X}}f(\mathbf{p},\mathbf{b},\mathbf{s})-\beta\cdot\text{cost}(\mathbf{p})$ (2)
For the search space $\mathcal{X}$, there could be implicit constraints during the configuration search. For example, the total number of GPUs must not exceed the available resources (e.g., 8 or 16). Alternatively, if the cloud service provider intends to fully utilize all available 8 GPUs, an implicit constraint could enforce the number of GPUs to be exactly 8. These constraints serve to reduce the search space. The system configuration involves parallelization configurations $\mathbf{p}$, maximum batch size configurations $\mathbf{b}$, and scheduling configurations $\mathbf{s}$. These variables are vectors, where each element corresponds to the configuration of an individual instance. Each instance is capable of handling requests independently, including managing (sub-)workers for tensor parallelism and pipeline parallelism. Instances within a stage process different requests in parallel, a concept referred to as data parallelism. Note that $\mathbf{p}$, $\mathbf{b}$, and $\mathbf{s}$ can have variable lengths, as the number of instances is itself a configurable parameter. For the $i$-th instance, we denote its stage as $\text{Stage}_{i}\in\{E,P,D\}$, where $E$ represents the encoder stage, $P$ represents the processing stage, and $D$ represents the decoder stage.
  • Parallelization: Let $\mathbf{p}$ denote the vector of parallel configs for all instances. For the $i$-th instance, if it is a prefill or decoding instance, then its config $p_{i}$ includes: $p^{\text{TP}}_{i}$, the number of GPUs used for tensor parallelism; and $p^{\text{PP}}_{i}$, the number of GPUs used for pipeline parallelism. If it is an encoding instance, considering IRP does not require communication, which is better than TP, we only use IRP. Therefore for encode, we overload the symbol $p^{\text{TP}}_{i}=p^{\text{IRP}}_{i}$ to denote the number of GPUs used for IRP. If the cost per GPU is a constant $c$, then the total cost is: $cost(\mathbf{p})=c\sum_{p_{i}\in\mathbf{p}}(p^{\text{TP}}_{i}\times p^{\text{PP}}_{i})$.
  • Scheduling involves two main decisions: First, which workers should each request be assigned to? Second, how should the order of requests be determined within a worker queue? To solve these decisions, we adapt strategies from the DistServe framework [21]. In the encoding stage, when a request arrives, it is assigned and pushed to an instance queue. Between different stages, global queues are used, and each available engine pulls proactively from the queues. Possible assignment strategies in the encoding stage include Round-Robin or Least-Loaded First for assigning requests. Once a request is assigned to a worker’s queue, we can apply ordering strategies like first-come-first-serve or shortest-job-first, or more complex strategies that prioritize requests based on their Service Level Objectives (SLOs). For simplicity, we constrain that all instances within the same stage share the same scheduling strategy.
Appendix B Implementation Details
B.1 Hyper-parameters in EPD system
We conducted our experiments using a cluster of 8 NVIDIA A100 GPUs (82GB). Each server was equipped with 128 CPUs and 1TB of RAM. The CUDA version was 12.2. Flash attention-2 was used for the attention implementation. We use FP16 precision for all experiments.
To ensure a fair comparison, we standardized key performance-affecting settings of the inference engine across all baselines. Specifically, these include a block size of 16; a maximum of 2048 blocks per request; context tokens capped at 49,152, and decoding tokens at 81,920 per batch. The scheduling policy for all stages was set to First-Come-First-Served (FCFS). Further, to allow enough resources for memory-heavy multimodal requests to execute, KV cache GPU utilization was set to 50%, and the maximum number of multimedia data of 32 was imposed per prompt. The size of the multimodal cache was fixed to 3000 across all models, and the vLLM inference engine was run in eager mode. Finally, for the vLLM inference engine, we used the version 0.6.1.post1 which represents a stable version for multimodal inference.
B.2 LMM Details
MiniCPM-V 2.6 integrates a SigLip-400M vision encoder, comprising 400 million parameters, with a Qwen2-7B language model, containing 7.6 billion parameters, culminating in 8 billion parameters. It excels in tasks involving single-image, multi-image, and video comprehension, even surpassing GPT-4V in these domains [19].
InternVL2-8B features an InternViT-300M-448px vision encoder with 300 million parameters, paired with an internlm2_5-7b-chat language model comprising 7.7 billion parameters, totaling 8 billion parameters. Similarly, InternVL2-26B combines an InternViT-6B-448px-V1-5 vision encoder with 6 billion parameters and an internlm2-chat-20b language model containing 20 billion parameters, resulting in 26 billion parameters.
B.3 SLO Criteria
Tab. 6 outlines the SLO criteria for TTFT and TPOT across various models and different numbers of images per request (#I/R). These criteria are empirically derived based on the characteristics of the underlying models, such as the computational complexity of the MME and LLM. We also consider what is realistically achievable by both our method and the baselines on a fixed number (8 GPUs) used in experiments. Across models, as #I/R increases, there is an approximately linear increase in the TTFT criteria due to the higher encoding load and additional tokens generated during the prefill stage. In contrast, TPOT requires only minor adjustments since it is not directly impacted by changes in #I/R.
MiniCPMv 2.6 8B InternVL 8B InternVL 26B
#I/R TTFT TPOT TTFT TPOT TTFT TPOT
2 1.40 0.04 1.20 0.05 3.50 0.07
4 2.60 0.04 2.40 0.06 7.05 0.08
6 3.90 0.06 3.55 0.09 11.00 0.95
8 5.10 0.06 5.00 0.18 15.00 0.15
Table 6: Requirements for TTFT and TPOT (in seconds) used for SLO computation across various models and number of images per request (#I/R) in the experiments.
B.4 Details for the Ablation Experiment of Offline Optimizer
In this experiment, 100 User requests arrive in real-time with each request containing 6 images, and they all request for the MiniCPMv model. For simplicity, we limit the configuration space to a restricted search space, where each data-parallel (DP) worker uses the same batch size configuration, and both TP and PP are fixed to 1. We explore the batch size configurations for workers across different stages, the number of DP workers in each stage, and the decision to enable IRP. The config identified by the optimizer is as follows: batch sizes for the E, P, and D stages are 2, 1, and 128, respectively; the number of workers in each stage is 6, 1, and 1; and IRP is enabled. We uniformly random-sample 10 configs from the search space. To maintain the same computational cost, all 8 GPUs are utilized across different configs. Consequently, the search space is constrained to leverage 8 GPUs, which can be enforced through rejection sampling. The sample mean of the performance metric (goodput, TTFT, and TPOT) is presented in Table 4. Recall that the goodput metric is defined as the maximum request rate while maintaining an SLO attainment of no less than 90%. For a fair comparison, When evaluating the TTFT and TPOT, we maintain the same request rate of 1.25 r/s, which is same as the goodput of the EPD system.
B.5 NPU Implementation Details
There are intrinsic differences between NPU and GPU architectures and processing characteristics. GPUs are general-purpose compute resources whereas NPUs have been specifically designed for neural networks. NPUs are optimized for FP16 calculations whereas GPUs are optimized for FP32 calculations. As stated above, CANN operators are used on Ascend NPU hardware in place of CUDA operators. These differences between CUDA and corresponding CANN operators, having been optimized for the hardware they are built upon, result in different latencies and trade-offs in inference as well as communication between NPU/GPU chips. This results in different encode/prefill/decode latencies which may be present different benefits from EPD and IRP.
The EPD-NPU system was implemented in a similar manner to our EPD-GPU system as described in Appendix B. All EPD-NPU experiments utilize IRP. The EPD-NPU distributed serving system was developed and deployed on a cluster of 910B3 NPUs, each with 64GB of high bandwidth memory. Some primary differences between the two implementations include:
  • Ascend-vLLM: EPD-NPU leverages Ascend-vLLM rather than vLLM which is an adaptation of vLLM on Ascend NPU hardware that utilizes the Compute Architecture for Neural Networks (CANN) software stack rather than CUDA.
  • Container-based deployment: Each encode, prefill, and decode instance is deployed separately as an API for the scheduler engine to call. These instances run in data parallel mode. The scheduler engine simply sends a POST request to the corresponding API with some minimal request-specific information (e.g., request ID, sampling parameters, block IDs to pull) and the remaining implementation mirrors that of EPD-GPU. This container-based deployment allows our framework to be easily scalable and portable to the cloud, providing intriguing future works regarding efficient utilization and scheduling of resources for users.
Figure 9: Left: The impact of tuning the number of encoding workers in the EPD method. The annotation ‘xEyP’ indicates that x encoders and y prefill workers are allocated for EPD, and ‘7P’ means that the DistServe always allocates 7 P workers for the encoding and prefill stage requests. Middle: The impact of the number of images per request in the workload. Right: The impact of encoding and prefill batch sizes in the EPD method (encoding and prefill batch sizes are set to be equal).
These differences in characteristics provide an opportunity to investigate the generalizability of our EPD framework and further understand the optimal setting/environment for disaggregating encode and prefill.
For EPD-NPU, we use eight 910B3 NPUs and evaluate the system by running the controlled experimental setup (as described in the experiments section of the main paper) in an online setting. Specifically, the number of input tokens was 22, length of output tokens was 10 with early stopping was disabled. The SLO requirements were also selected in a similar manner as the GPU experiments. The requirements for InternVL2-8B with 8 images per request were 8.5 and 0.12 for TTFT and TPOT, respectively. Each server had 192 CPUs and 1.5TB of RAM. The CANN version was 7.6. The maximum model length was set to 27000, KV Cache utilization was set to 83% And the maximum number of multimedia data of 8 was imposed. EPD-NPU also used a version of Ascend-vLLM (version 0.6.3.post1) built upon a newer version of vLLM than the one used in EPD-GPU. Unless otherwise stated, the hyper-parameters are identical to those described in B.1.
Appendix C Throughput in Offline Settings
In this section, we demonstrate the benefits of the EPD method in terms of throughput, specifically focusing on its memory efficiency and the ability to support higher batch sizes for both the E and P stages.
Consider an offline scenario where a batch of requests is submitted, allowing the system to process them overnight with a focus on maximizing end-to-end (E2E) throughput. This is an additional performance metric complementing the main body of the paper, measured in terms of requests completed per second and the system’s ability to handle high request volumes in offline settings. In the traditional DistServe method, memory is allocated for both the encoding and prefill stages on the same worker, which is constrained by the total memory capacity of the GPU.
This memory limitation becomes especially significant in heterogeneous environments with low-end GPUs that have limited memory. In these cases, the DistServe method struggles to fully utilize low-end GPUs, as the combined memory demands of the encoding and prefill stages exceed the available memory. In contrast, our disaggregated EPD method is more cost-effective, as it assigns the low-end GPUs exclusively to the encoding stage, allowing better utilization of available memory. In this experiment, we consider a less extreme scenario where the memory demands of the prefill workers in the DistServe method do not exceed the memory capacity of the low-end GPUs, allowing for a batch size of 1 for both the encoding and prefill stages.
(a) MiniCPMv-2.6                            (b) InternVL2-8B                        (c) InternVL2-26B
Figure 10: Columns (a), (b), and (c) correspond to MiniCPM-V 2.6, InternVL2-8B, and InternVL2-26B, respectively. The top and bottom rows show results for 6 and 8 images per request, respectively. As expected, EPD maintains a significant performance advantage over all baselines, demonstrating robustness of the system as the image count increases.
We use 8 A800 GPUs for this experiment. For EPD, we allocate 5 GPUs for encoding (E), 2 GPUs for prefill (P), and 1 GPU for decoding (D), with the maximum batch sizes set to 8, 8, and 128 for E, P, and D, respectively. In the DistServe method, we allocate 7 GPUs for prefill (P) and 1 GPU for decoding (D), with the maximum batch sizes of 1 and 128, respectively. The workload consists of 1,000 requests, each with a single image, a simple prompt (“What is the content of this image?”), and a maximum of 10 output tokens. We tune the system’s hyperparameters as outlined in Sec. 3.2.3 and evaluate their impact on throughput. The results are presented in Fig. 9.
The left plot illustrates the importance of selecting an appropriate GPU configuration. As described in Sec. 3.2.3, our algorithm automatically identifies that a 5E2P configuration (5 encoder and 2 prefill workers) maximizes E2E throughput. The middle plot shows that EPD achieves superior throughput when the number of images per request is low, demonstrating that EPD’s architecture prevents the encoder from becoming a compute-bound bottleneck. The right plot highlights that EPD’s performance is relatively insensitive to encoding and decoding batch sizes, consistently outperforming the DistServe method. This flexibility allows us to either choose a relatively large batch size or rely on the algorithm in Sec. 3.2.3 to automatically select an optimal batch size.
Appendix D Additional Experiments
D.1 SLO Attainment for End-to-End Generation for 6 and 8 images
In this section, we show additional results from the experiment in Sec. 4.1 pertaining to 6 and 8 images per request. As seen, the SLO attainment for EPD begins to decline as number of images rise, particularly at higher request rates. This decline reflects the increasing computational demand associated with processing multiple high-resolution images within a single request. Nevertheless, EPD continues to outperform all baselines, which struggle even at low request rates and fail to scale effectively under increased workloads.
@@ 40 OOM OOCL
Table 7: Comparison of maximum supported KV cache size (in terms of percentage of free memory) on prefill node for various #images/ request. Image resolution fixed to 4K. Higher (italicized) is better.
(a) GPU                        (b) NPU
Figure 11: Breaking down the latency of encode and prefill for InternVL2-8B across different number of images per requests. We compare the difference between GPU (left) and NPU (right) where light green is encode and light blue is prefill.
D.2 EPD Supports Larger KV Cache Sizes:
We compare the maximum KV cache size that can be allocated across systems. The batch size is fixed at 1, and the image resolution is set to 4032$\times$3024. As shown in Tab. 7, EPD supports significantly larger KV cache sizes than DistServe. For instance, for the InternVL2-26B model at 10 images per request, EPD can support a KV cache size of 80% vs 36% for DistServe, representing a 2.2$\times$ improvement. Notably, in certain scenarios—such as with the InternVL2-26B model and 20 images per request—the DistServe system encounters an OOM error, indicating that it cannot process 20 images even with a KV cache size of 0. Additionally, some configurations encounter an Out of Context Limit (OOCL) error, where the large number of encoding tokens generated by high image counts exceeds the LLM’s context limit during the prefill stage.
D.3 Heavier Encode in NPUs
We compare the encode-to-prefill stage latencies between NPUs and GPUs. Figure  11 show the results for the InternVL2-8B model, across various I/R values. As seen, we notice higher ($\approx 10-20\%$) encode-to-prefill latency ratios for NPUs in comparison to GPUs, indicating the need for a different optimized configuration for NPU.
This trend was consistent across different large multimodal models (LMMs), indicating that NPUs tend to spend a greater proportion of time in the encoding phase compared to GPUs.